For this problem, carry at least four digits after the decimal in your calculations. Answers may vary slightly due to rounding.

A random sample of 5913 physicians in Colorado showed that 3359 provided at least some charity care (i.e., treated poor people at no cost).

(a) Let p represent the proportion of all Colorado physicians who provide some charity care. Find a point estimate for p. (Round your answer to four decimal places.)


(b) Find a 99% confidence interval for p. (Round your answers to three decimal places.)

lower limit    
upper limit    

Give a brief explanation of the meaning of your answer in the context of this problem.

99% of the confidence intervals created using this method would include the true proportion of Colorado physicians providing at least some charity care.

(c) Is the normal approximation to the binomial justified in this problem? Explain.

Yes; np > 5 and nq > 5.
Answer #1

Solution :

Given that,

n = 5913

x = 3359

Point estimate = sample proportion = \hat p = x / n = 3359/5913=0.5681

1 - \hat p   = 1- 0.5681 =0.4319

At 99% confidence level the z is ,

\alpha  = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576 ( Using z table )

  Margin of error = E = Z\alpha/2   * (\sqrt(((\hat p * (1 - \hat p )) / n)

= 2.576* (\sqrt((0.5681*0.4319) / 5913)

E= 0.017

A 99% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.5681- 0.017< p <0.5681+ 0.017

0.551< p < 0.585

The 99% confidence interval for the population proportion p is : lower limit=0.551, upper limit=0.585